Stefan and Bill each improved their yards by planting hostas and shrubs. They bought their supplies from the same store. Stefan spent $117 on 7 hostas and 10 shrubs. Bill spent $108 on 8 hostas and 5 shrubs. What is the cost of one hosta and the cost of one shrub?
Answer by Boreal(15235)

x=hostas cost
y=shrubs cost
7x+10y=117
8x+5y=108
======
7x+10y=117
-16x-10y=-216, multiplying the bottom by (-2)
-9x=-99
x=11
substitute into first
77+10y=117
y=4
=====
check with second
88+20=108
hosts are $11
shrubs are $4